Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter whole numbers in the answer blank). Numerical Answers Expected


1
Expert's answer
2020-07-06T17:31:11-0400
"\textnormal{Let's imagine how a cone can be located}
\\
\textnormal{to occupy the maximum volume. }
\\
\textnormal{ Their bases and heights must be equal.}
\\
\textnormal{Let's write the known volume formulas}
\\
\textnormal{and calculate the answer}
\\
V_{cylinder} = S * h \implies h = \frac {V_{cylinder}} {S}
\\
V_{cone} = \frac{1}{3}*S*h=\frac{1}{3}*S*\frac {V_{cylinder}} {S}
\\
V_{cone} = \frac{1}{3}*\cancel{S}*\frac {V_{cylinder}} {\cancel{S}} = \frac{1}{3}*V_{cylinder}
\\
V_{cone} = \frac{1}{3}*15=5"